# Generative Bayesian optimization: generative models as acquisition functions

### v1

@@ 1 Introduction
Bayesian optimization (BO) has been a successful approach to solve complex black-box optimization problems by making use of probabilistic surrogate models, such as a Gaussian processes (GPs) (Rasmussen & Williams, 2006), and their uncertainty estimates (Shahriari et al., 2016; Garnett, 2023). BO methods have been particularly useful in areas such as hyper-parameter tuning for machine learning algorithms (Snoek et al., 2012), material design (Frazier & Wang, 2016), and robot locomotion (Calandra et al., 2016). The core idea of BO is to apply a Bayesian decision-theoretic framework to make optimal choices by maximizing an expected utility criterion, also known as an acquisition function. The corresponding expectations are taken under a Bayesian posterior over the underlying objective function. Thus, the Bayesian model provides a principled way to account for the uncertainty inherent to the limited amount of data and the noisy observations.
@@
In many applications such as simulated scenarios (Azimi et al., 2010), one is able to run multiple evaluations of the objective function in parallel, even though the simulations themselves might be expensive to run. Most common BO approaches to these batch settings incrementally build a set of candidates by sampling “fantasy” observations from the probabilistic model and conditioning on them before selecting the next candidate in the batch (Wilson et al., 2018). Although near-optimal batches can be selected this way, this approach is not scalable to very large batches in high-dimensional spaces, such as problems in protein design (Stanton et al., 2022; Gruver et al., 2023).
@@
One of the most promising alternatives to batch BO has been to train a generative model as a proposal distribution informed by the acquisition function and then sample a batch from the learned proposal (Brookes et al., 2019; Stanton et al., 2022; Gruver et al., 2023; Steinberg et al., 2025). This approach comes with several advantages. Firstly, given a trained generative model, sampling is usually inexpensive. Secondly, existing general-purpose generative models can be used and fine-tuned for the optimization task at hand. Lastly, sampling avoids estimating the global optimum of an acquisition function, which can be hard. However, existing generative approaches to black-box optimization usually rely on fitting a surrogate (regression or classification) model first then training a generative model on top of it (Stanton et al., 2022; Gruver et al., 2023; Steinberg et al., 2025). This two-stage process compounds approximation errors from both models and can increase the computational cost significantly when compared to having a single model.
@@
In this paper we present a general framework for learning generative models for batch Bayesian optimization tasks that requires a single model without the need for additional probabilistic regression or classification surrogates. Our approach encodes general utility functions into training objectives for generative models directly. We focus on two cases, one where we train the model via a loss function for a reward model analogously to the direct preference optimization (DPO) formulation for large language models (Rafailov et al., 2023), and the second one where we train the generative model through traditional divergence minimization, using utilities as part of sample weights. We present theoretical analyses of algorithmic convergence and performance and empirical results on practical applications involving high-dimensional combinatorial optimization problems.
@@ 2 Background
${\boldsymbol{\mathbf{x}}}^{*}\in\operatornamewithlimits{argmax}_{{\boldsymbol{\mathbf{x}}}\in{\mathcal{{X}}}}f({\boldsymbol{\mathbf{x}}})\,.$ (1)
The algorithm is restricted to a budget of $T\geq 1$ optimization rounds, where it chooses query locations ${\boldsymbol{\mathbf{x}}}_{t}\in{\mathcal{{X}}}$ to evaluate the objective function $f$. Observations are usually noisy $y_{t}=f({\boldsymbol{\mathbf{x}}}_{t})+\epsilon_{t}$, where $\epsilon_{t}$ is often assumed Gaussian or sub-Gaussian. Besides sequential decisions where a single ${\boldsymbol{\mathbf{x}}}_{t}$ is chosen per round, batch versions of BO take advantage of parallel function evaluations by selecting a batch of query points ${\mathcal{{B}}}_{t}:=\{{\boldsymbol{\mathbf{x}}}_{t,i}\}_{i=1}^{B}\subset{\mathcal{{X}}}$ per round. However, to simplify our presentation and avoid notation clutter, we will consider the single-point sequential setting to present the background and the initial derivation of our framework, later extending it to the batch setting.
@@
Assume a Gaussian process prior $f\sim\mathcal{GP}(0,k)$. Then, given a set of observations ${\mathcal{{D}}}_{t}:=\{{\boldsymbol{\mathbf{x}}}_{i},y_{i}\}_{i=1}^{t}$, with Gaussian noise $\epsilon\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$, the posterior $f|{\mathcal{{D}}}_{t}\sim\mathcal{GP}(\hat{f}_{t},k_{t})$ is available in closed form with mean and covariance function given by:
$\hat{f}_{t}({\boldsymbol{\mathbf{x}}})$ $:={\boldsymbol{\mathbf{k}}}_{t}({\boldsymbol{\mathbf{x}}})^{\mathsf{T}}({\boldsymbol{\mathbf{{K}}}}_{t}+\sigma_{\epsilon}^{2}{\boldsymbol{\mathbf{{I}}}})^{-1}{\boldsymbol{\mathbf{y}}}_{t}$ (2)
$k_{t}({\boldsymbol{\mathbf{x}}},{\boldsymbol{\mathbf{x}}}^{\prime})$ $:=k({\boldsymbol{\mathbf{x}}},{\boldsymbol{\mathbf{x}}}^{\prime})-{\boldsymbol{\mathbf{k}}}_{t}({\boldsymbol{\mathbf{x}}})^{\mathsf{T}}({\boldsymbol{\mathbf{{K}}}}_{t}+\sigma_{\epsilon}^{2}{\boldsymbol{\mathbf{{I}}}})^{-1}{\boldsymbol{\mathbf{k}}}_{t}({\boldsymbol{\mathbf{x}}}^{\prime})$ (3)
where ${\boldsymbol{\mathbf{k}}}_{t}({\boldsymbol{\mathbf{x}}}):=[k({\boldsymbol{\mathbf{x}}},{\boldsymbol{\mathbf{x}}}_{i})]_{i=1}^{t}\in\mathbb{{R}}^{t}$, ${\boldsymbol{\mathbf{{K}}}}_{t}:=[k({\boldsymbol{\mathbf{x}}}_{i},{\boldsymbol{\mathbf{x}}}_{j})]_{i,j=1}^{t}\in\mathbb{{R}}^{t\times t}$, ${\boldsymbol{\mathbf{y}}}_{t}:=[y_{i}]_{i=1}^{t}\in\mathbb{{R}}^{t}$, for ${\boldsymbol{\mathbf{x}}},{\boldsymbol{\mathbf{x}}}^{\prime}\in{\mathcal{{X}}}$. Having learned a model for $f$, one can compute an acquisition function $a_{t}({\boldsymbol{\mathbf{x}}})$ mapping candidate points ${\boldsymbol{\mathbf{x}}}\in{\mathcal{{X}}}$ to their expected utility $\mathbb{E}[u(y)|{\boldsymbol{\mathbf{x}}},{\mathcal{{D}}}_{t}]$, where the utility intuitively encodes how useful it is to collect an observation at $\boldsymbol{\mathbf{x}}$}. Classical examples of expected utility functions include the probability of improvement $a({\boldsymbol{\mathbf{x}}})=p(y\geq\tau|{\boldsymbol{\mathbf{x}}})=\mathbb{E}[\mathbb{I}[y\geq\tau]|{\boldsymbol{\mathbf{x}}}]$ and the expected improvement $a({\boldsymbol{\mathbf{x}}})=\mathbb{E}[\max\{y-\tau,0\}]$. The next candidate is then chosen as:
@@
When multiple evaluations can be executed in parallel, batch BO approaches become more appropriate. In this case, one selects a batch of candidates ${\mathcal{{B}}}_{t+1}=\{{\boldsymbol{\mathbf{x}}}_{t+1,i}\}_{i=1}^{B}\subset{\mathcal{{X}}}$ and runs multiple evaluations of the objective in parallel to collect a corresponding batch of observations. The difficulty there lies in ensuring the chosen batch of candidates is diverse enough (i.e., not having all candidates collapsed at the maximum of the acquisition function), while still ensuring optimality. To do so, batch BO methods often rely on a sequential greedy selection scheme and approximations thereof (Wilson et al., 2018; Garnett, 2023). Usually, the first batch point is selected by maximizing the acquisition function, then an observation is simulated by sampling from the surrogate model (e.g., a GP), which is used to condition the model as if it was an actual observation, and the process is repeated with the conditioned model until the given number $B$ of batch elements is selected. This process ensures a level of optimality to the chosen batch, but it is not easily scalable when very large batches are needed due to the cost of computing the simulated model updates. In contrast, one can train a generative model to produce samples guided by the GP surrogate in adaptive Bayesian quadrature (Kanagawa & Hennig, 2019) frameworks. For instance, $f$ can represent a log-likelihood or log-joint probability that depends on an expensive-to-evaluate computer model or real experimental process. We can then formulate an evidence lower bound with the GP surrogate model and train a generative model with it (Acerbi, 2020; Oliveira et al., 2021). Alternatively, in latent-space BO methods for high-dimensional problems (Gómez-Bombarelli et al., 2018; Stanton et al., 2022; Gruver et al., 2023), one learns a probabilistic representation of a (usually lower-dimensional) manifold of the data jointly with $f$, and performs BO in that space, projecting query points back to the original space at evaluation time. This technique has led to numerous BO methods for high-dimensional optimization (Gómez-Bombarelli et al., 2018; Gruver et al., 2023; González-Duque et al., 2024).
@@
Besides other approaches based on latent spaces, methods like variational search distributions (VSD, Steinberg et al., 2025) and batch BORE (Oliveira et al., 2022) learn a probabilistic classifier based on improvement labels and then match a generative model to the implied posterior. Namely, let $z_{t}:=\mathbb{I}[y_{t}>\tau]$, where $\tau$ is, e.g., the best observation so far or a (empirical) quantile of the marginal distribution of observations. We then learn a classification model $\pi$ for $p(z|{\boldsymbol{\mathbf{x}}})$ by minimizing a loss given by a proper scoring rule. For example, given the available data ${\mathcal{{D}}}_{t-1}$, the cross-entropy loss consists of:
$\ell_{t}[\pi]:=-\sum_{i=1}^{t-1}z_{i}\log\pi({\boldsymbol{\mathbf{x}}}_{i})+(1-z_{i})\log(1-\pi({\boldsymbol{\mathbf{x}}}_{i}))$ (5)
Given a prior $p_{0}$ over $\mathcal{{X}}$} and $\pi_{t}$ as the minimizer of $\ell_{t}$, we can now learn a generative model as:
@@
which corresponds to an evidence lower bound treating $\pi_{t}({\boldsymbol{\mathbf{x}}})\approx p(y>\tau|{\boldsymbol{\mathbf{x}}},{\mathcal{{D}}}_{t-1})$, so that the optimal variational distribution is $q_{t}({\boldsymbol{\mathbf{x}}})\approx p({\boldsymbol{\mathbf{x}}}|y>\tau,{\mathcal{{D}}}_{t-1})$.
@@
Large language models are typically fine tuned via reinforcement learning with human feedback (RLHF). This process involves training an LLM as a RL agent with a reward model. In practice, we do not directly observe rewards, but have access to user preferences. Given a prompt context $c$, let ${\boldsymbol{\mathbf{x}}}^{+},{\boldsymbol{\mathbf{x}}}^{-}\sim q({\boldsymbol{\mathbf{x}}}|c)$ denote two answers generated by an LLM $q$, with $\boldsymbol{\mathbf{x}}$$^{+}$ denoting the answer preferred by the user, and $\boldsymbol{\mathbf{x}}$$^{-}$ the dispreferred one. Given a dataset of user preferences ${\mathcal{{D}}}_{n}^{+}:=\{c_{i},{\boldsymbol{\mathbf{x}}}_{i}^{+},{\boldsymbol{\mathbf{x}}}_{i}^{-}\}_{i=1}^{n}$, one can then learn a reward model $\rho$ by minimizing the negative log-likelihood under a Bradley-Terry preference model:
$\ell_{n}^{+}[\rho]:=-\mathbb{E}_{(c,{\boldsymbol{\mathbf{x}}}^{+},{\boldsymbol{\mathbf{x}}}^{-})\sim{\mathcal{{D}}}_{n}^{+}}[\log\sigma(\rho(c,{\boldsymbol{\mathbf{x}}}^{+})-\rho(c,{\boldsymbol{\mathbf{x}}}^{-}))]$ (7)
Having learned a reward model $\rho_{n}$, RLHF trains the LLM as an agent. A KL-regularized objective reduces the risk of the optimized model deviating too much from a reference model $q_{\mathrm{ref}}$ given by a pre-trained LLM. The optimal generative model solves:
$q_{n}\in\operatornamewithlimits{argmax}_{q}\mathbb{E}_{c\sim{\mathcal{{D}}}_{n}^{+},{\boldsymbol{\mathbf{x}}}\sim q({\boldsymbol{\mathbf{x}}}|c)}[\rho_{n}(c,{\boldsymbol{\mathbf{x}}})]-\tau\mathbb{D}_{\mathrm{KL}}(q||q_{\mathrm{ref}})$ (8)
Direct preference optimization (DPO, Rafailov et al., 2023) removes the need for an explicit reward model by viewing the LLM itself through the lens of a reward model. It is not hard to show that, fixing a reward model $\rho$, the optimal solution to Equation 8 is given by:
$q({\boldsymbol{\mathbf{x}}}|c)=\frac{1}{\zeta_{\rho}(c)}q_{\mathrm{ref}}({\boldsymbol{\mathbf{x}}}|c)\exp\left(\frac{1}{\tau}\rho(c,{\boldsymbol{\mathbf{x}}})\right)\,$, (9)
where $\zeta_{\rho}(c):=\sum_{\boldsymbol{\mathbf{x}}}q_{\mathrm{ref}}({\boldsymbol{\mathbf{x}}}|c)\exp(\tau^{-1}\rho(c,{\boldsymbol{\mathbf{x}}}))$ is the partition function at the given context $c$. Although it would be intractable to evaluate $\zeta_{\rho}$ in practice, DPO uses the fact that, in the Bradley-Terry model, the partition function-dependent terms cancel out. Note that the reward model can be expressed in terms of $q$ as:
$\rho(c,{\boldsymbol{\mathbf{x}}})=\tau\log\left(\frac{q({\boldsymbol{\mathbf{x}}}|c)}{q_{\mathrm{ref}}({\boldsymbol{\mathbf{x}}}|c)}\right)+\tau\log\zeta_{\rho}(c)\,.$ (10)
@@
$\ell_{\mathrm{DPO}}[q]=-\mathbb{E}_{(c,{\boldsymbol{\mathbf{x}}}^{+},{\boldsymbol{\mathbf{x}}}^{-})\sim{\mathcal{{D}}}_{n}^{+}}\left[\log\sigma\left(\tau\log\left(\frac{q({\boldsymbol{\mathbf{x}}}^{+}|c)}{q_{\mathrm{ref}}({\boldsymbol{\mathbf{x}}}^{+}|c)}\right)-\tau\log\left(\frac{q({\boldsymbol{\mathbf{x}}}^{-}|c)}{q_{\mathrm{ref}}({\boldsymbol{\mathbf{x}}}^{-}|c)}\right)\right)\right]\,$, (11)
which eliminates the partition function $\zeta_{\rho}$. Therefore, we can train the generative model $q$ directly with $\ell_{\mathrm{DPO}}$ without the need for an intermediate reward model. Such simplification to a single training loop cuts down the need for computational resources, eliminates a source of approximation errors (from learning $\rho$), and brings in theoretical guarantees from Bradley-Terry models (Shah et al., 2016; Bong & Rinaldo, 2022). The main question guiding this work is whether we can apply a similar technique to simplify the training of (arbitrary) generative models for BO by removing the need for an intermediate classification or regression model.
Other related work.
There has been recent progress in adapting diffusion models for black-box optimization tasks, often done by learning a model that can be conditioned on observation values, learned from a dataset of evaluations (Krishnamoorthy et al., 2023). Other approaches involve guiding the diffusion process by a given utility function derived from a regression model (Gruver et al., 2023; Yun et al., 2025). Note, however, that such methodologies are specific to diffusion, whereas we focus on a general approach that can be applied to virtually any type of generative model.

### v2

@@ 1 Introduction
Bayesian optimization (BO) has been a successful approach to solve complex black-box optimization problems by making use of probabilistic surrogate models, such as a Gaussian processes (GPs) (Rasmussen2006), and their uncertainty estimates (Shahriari2016; Garnett2023book). BO methods have been particularly useful in areas such as hyper-parameter tuning for machine learning algorithms (Snoek2012), material design (Frazier2016), and robot locomotion (Calandra2016). The core idea of BO is to apply a Bayesian decision-theoretic framework to make optimal choices by maximizing an expected utility criterion, also known as an acquisition function. The corresponding expectations are taken under a Bayesian posterior over the underlying objective function. Thus, the Bayesian model provides a principled way to account for the uncertainty inherent to the limited amount of data and the noisy observations.
@@
In many applications such as simulated scenarios (Azimi2010), one is able to run multiple evaluations of the objective function in parallel, even though the simulations themselves might be expensive to run. Common BO approaches to these batch settings incrementally build a set of candidates by sampling “fantasy” observations from the probabilistic model and conditioning on them before selecting the next candidate in the batch (Wilson2018). Although near-optimal batches can be selected this way, this approach is not scalable to very large batches in high-dimensional spaces, such as problems in protein design (stanton2022accelerating; Gruver2023).
@@
One of the most promising alternatives to batch BO has been to train a generative model as a proposal distribution informed by the acquisition function and then sample a batch from the learned proposal (Brookes2019cbas; stanton2022accelerating; Gruver2023; Steinberg2025variational). This approach comes with several advantages. Firstly, given a trained generative model, sampling is usually inexpensive. Secondly, existing general-purpose generative models can be used and fine-tuned for the optimization task at hand. Lastly, sampling avoids estimating the global optimum of an acquisition function, which can be hard. However, existing generative approaches to black-box optimization usually rely on fitting a surrogate (regression or classification) model first then training a generative model on top of it (stanton2022accelerating; Gruver2023; Steinberg2025variational). This two-stage process compounds approximation errors from both models and can increase the computational cost significantly when compared to having a single model.
@@
In this paper we present a general framework for learning generative models for batch Bayesian optimization tasks that requires a single model without the need for additional probabilistic regression or classification surrogates. Our approach for generative BO (GenBO) encodes general utility functions into training objectives for generative models directly. We focus on two cases, one where we train the model via a loss function for a reward model analogously to the direct preference optimization (DPO) formulation for large language models (Rafailov2023dpo), and the second one where we train the generative model through divergence minimization, using utilities as part of sample weights. We present theoretical analyses on the convergence of approximations and empirical results on practical applications involving high-dimensional combinatorial optimization problems.
@@ 2 Background
${\boldsymbol{\mathbf{x}}}^{*}\in\operatornamewithlimits{argmax}_{{\boldsymbol{\mathbf{x}}}\in{\mathcal{{X}}}}f({\boldsymbol{\mathbf{x}}})\,$, (1)
where $f$ is an expensive-to-evaluate black-box function, i.e., $\nabla_{\boldsymbol{\mathbf{x}}}f$ is unavailable. We can only observe $f({\boldsymbol{\mathbf{x}}})$ via noisy evaluations $y=f({\boldsymbol{\mathbf{x}}})+\epsilon$, where $\epsilon$ is assumed sub-Gaussian (Pisier2016). We assume the objective $f$ can be evaluated in parallel, and the algorithm is allowed to run up to $T\geq 1$ optimization rounds with a batch of $B$ query locations ${\mathcal{{B}}}_{t}:=\{{\boldsymbol{\mathbf{x}}}_{t,i}\}_{i=1}^{B}\subset{\mathcal{{X}}}$ per round.
@@
Typically BO assumes a Bayesian prior over $f$ (Garnett2023book), often given by a Gaussian process (Rasmussen2006). Given a set of observations ${\mathcal{{D}}}_{t}$, corrupted by Gaussian noise $\epsilon\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$, the Bayesian posterior distribution over $f$ given the data ${\mathcal{{D}}}_{t}$ is available in closed form as a GP with known mean and covariance functions (see Appendix A). BO then uses the model’s posterior distribution to compute an acquisition function $a_{t}({\boldsymbol{\mathbf{x}}})$ mapping candidate points ${\boldsymbol{\mathbf{x}}}\in{\mathcal{{X}}}$ to their expected utility value $\mathbb{E}[u(y)|{\boldsymbol{\mathbf{x}}},{\mathcal{{D}}}_{t}]$, where the utility function $u$ intuitively encodes how useful it is to collect a new observation at $\boldsymbol{\mathbf{x}}$}. Classical examples of expected utilities include the probability of improvement $a_{t}({\boldsymbol{\mathbf{x}}}):=p(y\geq\tau|{\boldsymbol{\mathbf{x}}},{\mathcal{{D}}}_{t})=\mathbb{E}[\mathbb{I}[y\geq\tau]|{\boldsymbol{\mathbf{x}}},{\mathcal{{D}}}_{t}]$ and the expected improvement $a_{t}({\boldsymbol{\mathbf{x}}}):=\mathbb{E}[\max\{y-\tau,0\}|{\mathcal{{D}}}_{t}]$. The next candidate is then chosen as:
@@
This strategy can be extended to the batch setting in a variety of ways (Garnett2023book, §11.3). For instance, one can select the first batch point ${\boldsymbol{\mathbf{x}}}_{t,1}$ by maximizing $a_{t}$ as above, and then select the next candidate as ${\boldsymbol{\mathbf{x}}}_{t,2}\in\operatornamewithlimits{argmax}_{{\boldsymbol{\mathbf{x}}}\in{\mathcal{{X}}}}\mathbb{E}[u(y)|{\boldsymbol{\mathbf{x}}},{\mathcal{{D}}}_{t}\cup\{{\boldsymbol{\mathbf{x}}}_{t,1},\tilde{y}_{t,1}\}]$, where the expectation is over both $\tilde{y}_{t,1}\sim p(y|{\boldsymbol{\mathbf{x}}}_{t,1},{\mathcal{{D}}}_{t})$ and $y\sim p(y|{\boldsymbol{\mathbf{x}}},{\mathcal{{D}}}_{t}\cup\{{\boldsymbol{\mathbf{x}}}_{t,1},\tilde{y}_{t,1}\})$, and iterate over this process until $B$ candidates have been selected for parallel evaluation. Although near optimal, evaluating this conditional expectation becomes quickly intractable as the batch size grows. Hence, one usually resorts to Monte Carlo approximations (Wilson2018). Other BO strategies allow for efficient optimization of the batch in parallel, such as information-theoretic acquisition functions (Takeno2020; Teufel2024), or even asynchronously (Kandasamy2018). However, scaling up to large batches in high-dimensional domains, especially involving combinatorial or mixed discrete-continuous search spaces, remains challenging (Gonzalez-Duque2024).
@@
Instead of relying on a Bayesian surrogate model for $f$ and then computing an acquisition function $a$ on top of it, one can model $a$ directly, which is the main idea behind likelihood-free BO (Song2022lfbo). On this line, methods like variational search distributions (VSD, Steinberg2025variational) and batch BORE (Oliveira2022batch) learn a probabilistic classifier $\pi({\boldsymbol{\mathbf{x}}})\approx p(y\geq\tau)$ in the original space, $\mathcal{{X}}$}, based on improvement labels $z:=\mathbb{I}[y\geq\tau]$ and then generate batches by approximately sampling $B$ candidates from $p({\boldsymbol{\mathbf{x}}}|y\geq\tau)$. The classifier can be learned by, e.g., minimizing the cross-entropy loss:
$L_{n}(\pi):=-\sum_{i=1}^{n}z_{i}\log\pi({\boldsymbol{\mathbf{x}}}_{i})+(1-z_{i})\log(1-\pi({\boldsymbol{\mathbf{x}}}_{i}))\,.$ (3)
Given a prior $p_{0}$ over $\mathcal{{X}}$} and the classifier $\pi_{t}$ that minimizes $L_{n_{t}}$ over the current $n_{t}:=Bt$ data points in ${\mathcal{{D}}}_{t}$, we can now learn a generative model approximating $p({\boldsymbol{\mathbf{x}}}|y\geq\tau,{\mathcal{{D}}}_{t})$ as:
@@
which corresponds to an evidence lower bound treating $\pi_{t}({\boldsymbol{\mathbf{x}}})\approx p(y\geq\tau|{\boldsymbol{\mathbf{x}}},{\mathcal{{D}}}_{t})$ as a likelihood.
@@
The process above for learning $q_{t}$ can be likened to the typical fine-tuning of large language models (LLMs) via reinforcement learning with human feedback (RLHF, Bai2022rlhf), which would normally involve training the LLM as an RL agent with a reward model $\rho$. In practice, we do not directly observe rewards, but have access to user preferences. Given a prompt’s context $c$, corresponding to the RL state, let ${\boldsymbol{\mathbf{x}}}^{+},{\boldsymbol{\mathbf{x}}}^{-}\sim q({\boldsymbol{\mathbf{x}}}|c)$ denote two answers generated by an LLM $q$, with $\boldsymbol{\mathbf{x}}$$^{+}$ denoting the answer preferred by the user, and $\boldsymbol{\mathbf{x}}$$^{-}$ the dispreferred one. Having a dataset ${\mathcal{{D}}}_{n}^{+}:=\{c_{i},{\boldsymbol{\mathbf{x}}}_{i}^{+},{\boldsymbol{\mathbf{x}}}_{i}^{-}\}_{i=1}^{n}$, one can then learn a reward function $\rho$ by minimizing the negative log-likelihood under a preference model, such as Bradley1952rank:
$L_{n}^{+}(\rho):=-\mathbb{E}_{(c,{\boldsymbol{\mathbf{x}}}^{+},{\boldsymbol{\mathbf{x}}}^{-})\sim{\mathcal{{D}}}_{n}^{+}}[\log\sigma(\rho(c,{\boldsymbol{\mathbf{x}}}^{+})-\rho(c,{\boldsymbol{\mathbf{x}}}^{-}))]\,.$ (5)
Having learned a reward model $\rho_{n}$, RLHF trains the LLM as to approximate an agent’s optimal policy under $\rho_{n}$. Regularization based on the Kullback-Leibler (KL) divergence with respect to a reference model $q_{\mathrm{ref}}$ is further added to improve stability. The optimal generative model then solves:
$q_{n}\in\operatornamewithlimits{argmax}_{q}\mathbb{E}_{c\sim{\mathcal{{D}}}_{n}^{+},{\boldsymbol{\mathbf{x}}}\sim q({\boldsymbol{\mathbf{x}}}|c)}[\rho_{n}(c,{\boldsymbol{\mathbf{x}}})]-\beta\mathbb{D}_{\mathrm{KL}}(q||q_{\mathrm{ref}})\,.$ (6)
Direct preference optimization (DPO, Rafailov2023dpo) removes the need for an explicit reward model by viewing the LLM itself through the lens of a reward model. It is not hard to show that, fixing a reward model $\rho$, the optimal solution to Equation 6 is given by:
$q({\boldsymbol{\mathbf{x}}}|c)=\frac{1}{\zeta_{\rho}(c)}q_{\mathrm{ref}}({\boldsymbol{\mathbf{x}}}|c)\exp\left(\frac{1}{\beta}\rho(c,{\boldsymbol{\mathbf{x}}})\right),$ (7)
where $\zeta_{\rho}(c):=\sum_{\boldsymbol{\mathbf{x}}}q_{\mathrm{ref}}({\boldsymbol{\mathbf{x}}}|c)\exp(\beta^{-1}\rho(c,{\boldsymbol{\mathbf{x}}}))$ is the partition function at the given context $c$. Although it is intractable to evaluate $\zeta_{\rho}$ in practice, DPO uses the fact that, in the Bradley-Terry model, the partition function-dependent terms cancel out. Note that the reward model $\rho$ can be expressed in terms of the optimal $q$ as:
$\rho(c,{\boldsymbol{\mathbf{x}}})=\beta\log\left(\frac{q({\boldsymbol{\mathbf{x}}}|c)}{q_{\mathrm{ref}}({\boldsymbol{\mathbf{x}}}|c)}\right)+\beta\log\zeta_{\rho}(c)\,.$ (8)
@@
$L_{\mathrm{DPO}}(q)=-\mathbb{E}_{(c,{\boldsymbol{\mathbf{x}}}^{+},{\boldsymbol{\mathbf{x}}}^{-})\sim{\mathcal{{D}}}_{n}^{+}}\left[\log\sigma\left(\beta\log\left(\frac{q({\boldsymbol{\mathbf{x}}}^{+}|c)}{q_{\mathrm{ref}}({\boldsymbol{\mathbf{x}}}^{+}|c)}\right)-\beta\log\left(\frac{q({\boldsymbol{\mathbf{x}}}^{-}|c)}{q_{\mathrm{ref}}({\boldsymbol{\mathbf{x}}}^{-}|c)}\right)\right)\right]\,$, (9)
which eliminates the partition function $\zeta_{\rho}$ terms. Therefore, we can train the generative model $q$ directly with $L_{\mathrm{DPO}}$ without the need for an intermediate reward model. Such simplification to a single training loop cuts down the need for computational resources, eliminates a source of approximation errors (from learning $\rho$), and brings in theoretical guarantees from Bradley-Terry models (Shah2016; Bong2022). The main question guiding our work is whether we can apply a similar technique to simplify the training of (arbitrary, not necessarily LLM) generative models for likelihood-free BO by removing the need for an intermediate surrogate model for $f$.